A carpenter has a 14 in piece of molding. After he cuts a 8 1/3 in piece from this piece of molding, how long is the original piece of molding, How long is the original piece of molding?
Answer by checkley77(12844)

14-8 1/3=42/3-25/3=17/3 or 5 2/3 inches remaing.
14+8 1/3=42/3+25/3=67/3 or 22 1/3 inches for the original piece.